Database last updated on Sun, 01 Feb 2026 19:40:55 +0000 by localhost
Instrument constant changes for the week before 2023307 (2 sec, 0 rows)
HG burned out? (AP results when HG is ON) (0 sec, 17 rows)
Measurements and tests summary.
_jday
_loc
_serial
_skc
dt
ed
fm
fz
hg
hp
rs
sl
uv
307
RESOLUTE
#205
ecsc,Next schedule will be ecsc (now ecsc ZS= 93 ZS= 93.075)
1
2
53
4
10
10
1
8
35
(0 sec)
Dead time (0 sec)
File name
Count
Serial
Expected (+/-5)
Measured
30-day history
B30723.205
1
#205
30
27